4. Consider the following contingency table:
Under 20 21-30 31-40
Male 12 12 17
Female 13 16 21

a. If one person is selected at random, what is the probability that person is Female? ______
b. If one person is selected at random, what is the probability that person is either under 20 or over 30? _______
c. If one person is selected at random, what is the probability that person is either male or in the age group 21-30? _____
d. If two people are selected at random, what is the probability they are both female? ______

Answer by oscargut(2103)

Total are 91 persons, 41 Male and 50 Female
Under 20 are 25 persons, over 30 are 38 persons
Male or in he group 21-30 are 41+16=57
a. If one person is selected at random, what is the probability that person is Female? _50/91_____
b. If one person is selected at random, what is the probability that person is either under 20 or over 30? ___63/91____
c. If one person is selected at random, what is the probability that person is either male or in the age group 21-30? _57/91____
d. If two people are selected at random, what is the probability they are both female? ___(50/91)(50/91)=2500/8281___